A motor pumps 50 kg of water into a tower which is 35 m high. It does this in a
time of 5 minutes. What is the power of the motor in this case?
1
Expert's answer
2020-11-25T10:53:28-0500

We can find the power of the motor from the definition of the power:


"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{50\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot 35\ m}{5\cdot 60\ s}=57\ W."

Answer:

"P=57\ W."